Instruments & Operations



Older News and Updates

06 April 2012

LUCI is currently undergoing testing in the instrument lab. We are still on schedule to restart normal science operations for the April science block. Here we provide some details about the change to the new detector and other things that PIs need to be aware of when planning their observations. We will update this page as the re-calibration proceeds, so please check back here often!

New detector

Other Instrument Changes

Exposure Time Calculator


Script Checker (v4.21, 06.April.2012)

Example data from the new detector:

Raw Twilight Flatfield Data in the Archive:



Recommendations to all PIs planning observations starting March 2012

May 2011: The LUCI Team has traced the likely source of the stray light seen in low-background observations to be caused by the shield temperature stabilization system sitting behind the detector. One of the heater resistors appears to be running hotter than usual, emitting thermal IR photons that reflect off the filter and down to the detector. The level is at a low surface brightness (~6 DN/pixel in 300s) and reasonably stable. An example image can be found here.

This stray light is effectively excess dark current and will largely subtract out when you sky subtract your data. Observers should make sure properly dithered data are taken to allow for a good subtraction of this excess background, although this is the usual way IR data are taken. The largest impact will be seen in data with the lowest background and longest exposure times, such as short wavelength (z, Y, or J band) spectroscopy and/or narrowband imaging.

Further investigation and repair of this problem scheduled during the normal summer shutdown and maintenance of LUCI1.

May 2011: LUCI script syntax checker updated to version 4.6, now including a visualization of where the dither positions will land with respect to the LUCI field of view. Go to this link for more information.

February 2010: LUCI script syntax checker updated to version 2.3. This fixes patrol field issues at high declinations.

January 2010: LUCI script syntax checker released

Steve Allanson's LUCI script syntax checker has been released today (21 Jan 2011) for use by the partnership. Go to this link for more information.

January 2010: MOS unit operable after a software patch

In December one of the two sensors that detects when a mask is in position at the focal plane unit apparently failed. A software patch in January has restored functionality and the MOS unit can again be used for science.

October 2010: Flexure Compensation Issue

Effective immediately: ALL scripts should have FLEXURE_COMP = OFF

Repeated tests with LUCIFER are showing that the shape of the flexure compensation curves are continuously changing while the maximum amplitude of 12pix/6pix (N3.75/1.80) is not increasing. The result is that the image motion on the detector with flexure compensation can partially be worse than without any compensation. Therefore, we recommend that observations be done with flexure compensation OFF until further notice.

Note that the flexure compensation NEVER took place during an integration in order to avoid possible increased detector noise/vibrations from driving the stepper motors. Thus there is NO decreased performance for single

Without flexure compensation the shift in between the individual exposures will be increased with regard to former observations, both in the spectral dispersion direction and in the spatial direction. The maximum/mean shift is 0.7/0.4 pixels per 15 degree change of rotator angle. This translates to about 1 hour total observing at EL=45deg, 30 minutes at EL=65deg and 15 minutes at EL=75deg.

The consequences are:

October 2010: Changes to the permanent slits

As part of the maintenance run during the 2010 summer shutdown, two new longslits have been installed: 2.0 and 1.5 arcsec wide. Also, the positions of the permanent slits have been moved. If your observing scripts referred to masks by their cabinet number (NB#) then you need to update the scripts to reflect the current mask positions. See the Observer's Reference Page on the wiki under the LUCIFER Installed MOS Masks link for the current mask positions. As always, you can refer to the mask ID numbers (e.g. ID990034 is the 1.0 arcsec slit) as these are cabinet position independent.

July 2010: The LUCIFER Team completed a maintenance run during the summer shutdown, including:

July 2010: the LUCIFER Mask Software (LMS) has been updated to version 1.60 (lms_v160.tar.gz, 10 MB). LMS runs on Macs with minimal configuration changes, see these instructions for installation.